# jina-embeddings-v3: Multilingual Embeddings With Task LoRA

### v1

@@ 4.3.1 Classification Adapter
From each dataset, we construct tuples consisting of two text values from the same class $(q,p)$ and seven text values from different classes $(n_{1},\dots,n_{7})$, resulting in a tuple of nine text values $(q,p,n_{1},\dots,n_{7})$. The model is trained to assign a high cosine similarity to the embeddings of $q$ and $p$, while enforcing low cosine similarity between $(q,n_{i})$. Each batch is composed of tuples sampled from a single dataset.
An extended version of the InfoNCE loss $\mathcal{L}_{\text{triplet}}$ from our previous work incorporates these additional negative samples:
@@
When using this loss function, texts from the same class as text $q_{i}$ that occur in the same batch as negatives for a different text $q_{j}$ $(i\neq j)$ are also treated as negatives. This introduces the issue of false negatives.
@@ 4.3.2 Text Matching Adapter
This adapter is trained to produce embeddings that quantify the similarity between two text values. It is applicable for tasks such as semantic textual similarity (STS) and retrieval tasks where there is a clear distinction between query and target text values. An example of such a retrieval task is duplicate detection, where text values from a corpus are compared against each other. In these cases, the “query” and “corpus” texts are treated symmetrically.
@@
To train the model with this objective, we utilize data from semantic textual similarity (STS) datasets such as STS12 Agirre et al. (2012) and SICK Marelli et al. (2014). These datasets consist of triplets $(q_{i},p_{i},t_{i})\in D$, where $(q_{i},p_{i})$ are text pairs and $t_{i}$ is the corresponding relevance score. A batch $B$ is constructed by selecting text values from a given number of triplets, with the ground truth similarity defined as $\zeta(q_{i},p_{i})=t_{i}$.
To enhance the model’s performance across languages, we translate the STS12 and SICK datasets into multiple languages using machine translation models such as WMT19 Ng et al. (2019) and MADLAD-3B Kudugunta et al. (2023). While training on STS datasets is highly effective, obtaining large volumes of this data is challenging due to the human annotation required. As a result, we incorporate various natural language inference (NLI) datasets into the training process.
Following previous work Wang et al. (2022), we prepare paired datasets with multiple negatives and apply the $\mathcal{L}_{\mathrm{triplet}}$ loss function. During each training step, either an STS or NLI dataset is selected, and the batch is constructed solely from the chosen dataset, using the appropriate loss function. In other words, each batch contains text values from only one specific dataset.
The relevant hyperparameters are provided in Appendix A1.
@@ 4.3.3 Asymmetric Retrieval Adapters
As discussed in Section 2.3, asymmetric retrieval tasks, such as question answering and traditional information retrieval, perform better with distinct encoding mechanisms for queries and documents. In this work, we follow the method proposed by Wang et al. (2022), using two distinct prefixes, but further separate the encoding processes by employing two specialized adapters, which are trained jointly. A detailed ablation study demonstrating the effectiveness of this approach is presented in Section 5.5.2.
@@ 4.3.5 Separation Adapter
The separation adapter is designed to perform well on clustering and reranking tasks. It is trained to distinguish between text values belonging to the same group and those from different groups. For reranking tasks, the adapter separates relevant from irrelevant documents based on a query’s information need. In clustering tasks, groups of text values are provided, and after calculating the embeddings and applying a clustering algorithm (e.g., k-means), the resulting clusters should reflect the correct groupings.
@@ 5.2 Performance on MTEB
Consulting the English MTEB leaderboard, it is noteworthy that embedding models built on LLMs perform only marginally better than jina-embeddings-v3, but come with increased complexity in real-world applications. For instance, e5-mistral-7b-instruct achieves an average score of 66.63 across all 56 English MTEB tasks (approximately 1.03% higher than jina-embeddings-v3), but produces embeddings with a dimension of 4096 and has a parameter size of 7.1 billion. In contrast, jina-embeddings-v3 has an embedding dimension of 1024 (with the option to reduce dimensions using MRL with a small performance trade-off, as discussed in Section 5.5.1) and only 570 million parameters, making it far more practical for downstream applications.
@@ 5.3 Performance on LongEmbed MTEB
We evaluate our model against text-embedding-3-large, bge-m3, and our previously released model suite jina-embeddings-v2 on six long document retrieval tasks from the MTEB leaderboard. The results, presented in Table 4, demonstrate that jina-embeddings-v3 with the text-matching adapter achieves the highest average performance. These findings underscore the effectiveness of the RoPE-based positional embeddings, outperforming both the fixed positional embeddings used by bge-m3 and the ALiBi-based approach employed in jina-embeddings-v2.
@@ 6 Conclusion
In this paper, we introduced jina-embeddings-v3, our latest text embedding model. By leveraging task-specific adapter tuning and failure-aware synthetic data augmentation on top of a robust backbone, jina-embeddings-v3 demonstrates competitive performance across a wide range of tasks. Extensive evaluations on both English and multilingual datasets highlight the model’s strong performance while maintaining a reasonable parameter size.
For future work, we plan to focus on enhancing the model’s performance for low-resource languages, further strengthening its capabilities in multilingual tasks where data availability is limited.

### v2

@@ 4.3.1 Classification Adapter
From each dataset, we construct tuples consisting of two text values from the same class $(q,p)$ and seven text values from different classes $(n_{1},\dots,n_{7})$, resulting in a tuple of nine text values $(q,p,n_{1},\dots,n_{7})$. The model is trained to assign a high cosine similarity to the embeddings of $q$ and $p$, while enforcing low cosine similarity between $q$ and all $n_{i}$. Each batch is composed of tuples sampled from a single dataset.
We employ an extended version of the InfoNCE loss $\mathcal{L}_{\text{triplet}}$ described in our previous work (CITATION HERE) to take into account these additional negative samples.
@@
When using this loss function, texts from the same class as text $q_{i}$ that occur in the same batch as negatives for a different text $q_{j}$ $(i\neq j)$ are also treated as negatives. This introduces the problem of false negatives.
@@ 4.3.2 Text Matching Adapter
This adapter is trained to produce embeddings that quantify the similarity between two text values. It is applicable for tasks such as semantic textual similarity (STS) and retrieval tasks where there is no clear distinction between query and target text values. An example of such a retrieval task is duplicate detection, where text values from a corpus are compared against each other. In these cases, the “query” and “corpus” texts are treated symmetrically.
@@
To train the model with this objective, we use data from semantic textual similarity (STS) training datasets such as STS12 Agirre et al. (2012) and SICK Marelli et al. (2014). These datasets consist of triplets $(q_{i},p_{i},t_{i})\in D$, where $(q_{i},p_{i})$ are text pairs and $t_{i}$ is the corresponding relevance score. A batch $B$ is constructed by selecting text values from a given number of triplets, with the ground truth similarity defined as $\zeta(q_{i},p_{i})=t_{i}$.
To enhance the model’s performance across languages, we translate the STS12 and SICK datasets into multiple languages using machine translation models, i.e. WMT19 Ng et al. (2019) and MADLAD-3B Kudugunta et al. (2023). While training on STS datasets is highly effective, obtaining large volumes of this data is challenging due to the human annotation required. As a result, we have incorporated various natural language inference (NLI) datasets into the training process.
During each training step, either an STS or NLI dataset is selected, and the batch is constructed solely from the chosen dataset, using the appropriate loss function. In other words, each batch contains text values from only one specific dataset.
For the relevant hyperparameters, see Appendix A1.
@@ 4.3.3 Asymmetric Retrieval Adapters
As discussed in Section 2.3, asymmetric retrieval tasks, such as question answering and traditional information retrieval, perform better with separate encoding mechanisms for queries and documents. In this work, we follow the method proposed by Wang et al. (2022), using two distinct prefixes, but further separate the encoding processes by employing two specialized adapters, which are trained jointly. A detailed ablation study demonstrating the effectiveness of this approach is presented in Section 5.5.2.
@@ 4.3.5 Separation Adapter
The separation adapter is designed to perform well on clustering and reranking tasks. It is trained to distinguish between text values belonging to the same group and those from different groups. For reranking tasks, the adapter separates relevant from irrelevant documents based on a query’s information need. In clustering tasks, groups of texts are provided, and after calculating the embeddings and applying a clustering algorithm (e.g., k-means), the resulting clusters should reflect the correct groupings.
@@ 5.2 Performance on MTEB
Consulting the English MTEB leaderboard, we plotted the performance of the top 100 embedding models against their parameter sizes in Figure 7. Notably, embedding models built on large language models (LLMs) perform only marginally better than jina-embeddings-v3 but entail significantly higher complexity and computational costs in real-world applications. For example, e5-mistral-7b-instruct achieves an average score of 66.63 across all 56 English MTEB tasks (approximately 1.03% higher than jina-embeddings-v3), but produces embeddings with a dimension of 4096 and has a parameter size of 7.1 billion. In contrast, jina-embeddings-v3 offers a more practical solution, with an embedding dimension of 1024 (or lower, using MRL with a manageable performance trade-off, as discussed in Section 5.5.1) and only 570 million parameters.
@@ 5.3 Performance on LongEmbed MTEB
We have evaluated our model against text-embedding-3-large, bge-m3, and our previously released model suite jina-embeddings-v2 on six long document retrieval tasks from the MTEB leaderboard. The results, presented in Table 4, demonstrate that jina-embeddings-v3 with the text-matching adapter achieves the highest average performance. These findings underscore the effectiveness of the RoPE-based positional embeddings, outperforming both the fixed positional embeddings used by bge-m3 and the ALiBi-based approach employed in jina-embeddings-v2.
@@ 6 Conclusion
This paper introduces jina-embeddings-v3, our latest text embedding model. By leveraging task-specific adapter tuning and failure-aware synthetic data augmentation on top of a robust backbone, jina-embeddings-v3 demonstrates competitive performance across a wide range of tasks. Extensive evaluations on both English and multilingual datasets highlight the model’s strong performance while maintaining a reasonable parameter size.
We are particularly interested in evaluating and improving the model’s performance on low-resource language and further analyzing systematic failures caused by low data availability. We plan to focus on this area going forward, further strengthening its capabilities in multilingual tasks where data availability is limited.
@@ Appendix A Appendix
Figure 2: Scaling law of embedding models. Each dot represents an embedding model. jina-embeddings-v3 demonstrates superior performance compared to models of similar size, showing a superlinear improvement over its predecessor, jina-embeddings-v2. This graph was created by selecting 100 embedding models from the MTEB leaderboard [^7] https://huggingface.co/spaces/mteb/leaderboard, excluding those without size information, typically closed-source or proprietary models. Submissions identified as outliers or trolling were also filtered out. The average MTEB performance on English tasks is plotted against the number of model parameters. The trendline, representing all models, is highlighted, with multilingual models emphasized in orange.